OpenStudy (anonymous):

p+9q+9+14p 6b++6b2+4b^3, and a^2+2b+2a^2+b+2c ... help? I don't understand, show work. Please? O_o I REALLY don't get it... thanks...=)

OpenStudy (anonymous):

Simplify.

OpenStudy (anonymous):

Thanks to everyone!!!!!! :)

OpenStudy (anonymous):

p+9q+9+14p = rewrite p+14p+9q+9= add like terms 15p+9q+9= factor out 3 3(5p+3q+3)

OpenStudy (anonymous):

6b+6b^2+4b^3=all are even => factor out 2 and b 2b(3+3b+2b^2)

OpenStudy (anonymous):

a^2+2b+2a^2+b+2c = regrouping a^2+2a^2+2b+b+2c = like terms 3a^2+3b+2c = could go father, but not necessary 3(a^2+b) +2c